The amount of an investment of P dollars for t years at simple interest rate r is given by A = P + Prt.
a) Rewrite this formula by factoring out the greatest
common factor on the right-hand side.
b) Find A if $8300 is invested for 3 years at a simple
interest rate of 15%.

Answer by ankor@dixie-net.com(22740)

The amount of an investment of P dollars for t years at simple interest rate r
is given by A = P + Prt.
:
a) Rewrite this formula by factoring out the greatest
common factor on the right-hand side.
A = P(1 + rt)
;
b) Find A if $8300 is invested for 3 years at a simple
interest rate of 15%.
A = 8300(1 + (.15*3))
A = 8300 * 1.45
A = $12,035